Two semiconductor manufacturing companies decide to merge in order to reduce cost and improve productivity. Each company has high speed test machines as well as other test machines. One company has 400 chip test machines of which 20% are high speed. The combined company will have 44% high speed test machines. If the second company has 60% high speed test machines, then how many machines are there in the second company in total?

Answer by 303795(562)

The first company has 20/100*400=80 high speed machines.
Let company 2 have x machines.
The combined number of machines is 400 +x.
Company 2 has 60/100x = 0.6x high speed machines.
The total number of high speed machines=80 + 0.6x
so high speed machines = 44/100 x (total machines)
80 + 0.6x = 0.44(400 + x)
80 + 0.6x = 176 + 0.44x
0.16x = 96
x=96/0.16 = 600
So the second company has 600 machines (360 of which are high speed)